# references.bib

### v2

@@ 5.2 Effect of Parameters $s$ and $t$
In the BCDNS method, the parameters $s$ and $t$ play a crucial role. Here, we vary the parameters $s$ and $t$ to solve Test Problem 1. The number of reduced cost computations serves as a key indicator in the BCDNS method. For each case, we record the corresponding number of pivots and reduced cost computations. And then we calculate the averages of these metrics to evaluate the algorithm’s performance. In this context, the number of pivots refers to the number of iterations performed by the NS method.
Figure 2: Effects of $s$ and $t$ on the number of pivots and reduced cost computations
Figure 1 illustrates the effects of $s$ and $t$ on two key performance metrics: the number of pivots and the total number of reduced cost computations. We can observe that the number of pivots decreases as $s$ increases, as larger subproblem sizes reduce the total iterations required by the NS method. However, as $s$ increases, the size of each subproblem also grows, which, while reducing the number of pivots, increases the computational cost of each pivot. Similarly, increasing $t$ enhances the quality of the variables selected for the subproblem, further reducing the number of pivots.
For small values of $s$, increasing $s$ reduces the total number of reduced cost computations by enabling more efficient solutions to larger subproblems. However, beyond a certain threshold, further increases in $s$ lead to an overall rise in computational effort due to the greater complexity of individual subproblems. The effect of $t$ is less pronounced, except when $s$ is very small, in which case increasing $t$ significantly increases the computational burden, as the algorithm approaches the behavior of the standard NS method. Extensive numerical experiments consistently reveal that reduced cost computations are minimized when $t$ is approximately $2s$.
@@ 5.3 Comparison with the Existing Methods
In this chapter, we primarily evaluate the performance differences among the GS-BCDNS method, the Sinkhorn algorithm, and the NS method.
Figure 3: Comparison of algorithms to computer Wasserstein distance
As described in Section 4.1, we apply the Sinkhorn algorithm with different values of $\epsilon$ and the GS-BCDNS method to solve Test Problems 2, 3 and 4. The stopping criterion for the Sinkhorn algorithm is set such that the algorithm terminates when the change in the objective function value falls below $\delta=10^{-6}$ or when the runtime exceeds 2000 seconds. The GS-BCDNS method is tested with fixed parameters $s=5/n$ and $t=10/n$. In the evaluation, the x-axis represents the runtime, while the y-axis represents the gap between the current objective function value and the optimal objective function value, where the optimal solution is obtained using the interior-point method implemented in Python’s SciPy library.
The results in Figure 3 indicate that for larger values of $\epsilon$ (i.e, $\epsilon=0.1$ and $\epsilon=0.01$), the Sinkhorn algorithm exhibits a faster convergence speed than the GS-BCDNS method; however, the approximation error remains relatively large. When $\epsilon=0.001$, the convergence speed of the Sinkhorn algorithm is comparable to that of the GS-BCDNS method, but its final convergence is slightly slower. For extremely small values of $\epsilon$, the convergence speed of the Sinkhorn algorithm is significantly slower than that of the GS-BCDNS method.
Figure 4: Comparison of algorithms to solve Test Problem 5
Figure 4 presents the results of solving Test Problem 5 using both the NS method and the GS-BCDNS method. The y-axis of Figure 4 represents the ratio of the NS method’s runtime to that of the GS-BCDNS method. The results demonstrate that the GS-BCDNS method significantly outperforms the NS method across all tested datasets.
@@ 6.   Conclusion
In this paper, we review the existing methods for solving the OT problems, including the NS method, the BCD method, and the Sinkhorn algorithm. Then we propose the BCDNS methods and explain thier theoretical foundations and the basis variable succession strategy. Furthermore, we provide a detailed algorithmic framework and established theoretical guarantees for finite termination and convergence to the exact optimal solution.
Extensive numerical experiments demonstrate that the BCDNS method achieves superior performance in terms of runtime and computational cost compared to the NS method. While the BCDNS method converges slightly slower than the Sinkhorn algorithm with large values of $\epsilon$, it has the advantage of guaranteeing convergence to an exact solution. Moreover, compared to the Sinkhorn algorithm with small values of $\epsilon$, the BCDNS method exhibits even faster convergence. This makes the BCDNS method particularly valuable for large-scale OT problems that require exact solutions.
Future work could explore several directions. First, improving the selection strategy for block coordinates may further enhance convergence speed. Second, extending the BCDNS method to more general network flow problems could broaden its applicability.
Appendix A Proof of $\operatorname{rank}(A)=M-1$
Proof.
Let
$A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -\mathbf{1}_{n}^{\top}&0&\cdots&0\\ 0&-\mathbf{1}_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\mathbf{1}_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},\quad M=n+m,\;N=nm,$
where $I_{n}$ is the $n\times n$ identity matrix and $\mathbf{1}_{n}$ is the $n$-vector of all ones.
(i) $\operatorname{rank}(A)\leq M-1$.
Denote the rows of $A$ by $r_{1},\dots,r_{M}$. Observe that
[…]
because each of the first $n$ rows contributes exactly one “$+1$” in each block and each of the last $m$ rows contributes exactly one “$-1$”, so they cancel out. Hence the rows are linearly dependent, and
$\operatorname{rank}(A)\;\leq\;M-1.$
(ii) $\operatorname{rank}(A)\geq M-1$.
Interpret $A$ as the node–arc incidence matrix of the complete bipartite directed graph $K_{n,m}$ with $M=n+m$ nodes and $nm$ arcs. Choose any spanning tree $T$ of this graph, which has exactly $M-1$ edges. Let
$B\;=\;A(:,E(T))\;\in\;\mathbb{R}^{M\times(M-1)}$
be the submatrix of $A$ whose columns correspond to the edges in $T$. Since $T$ is a tree, its incidence matrix has rank $M-1$, and in particular there is an $(M-1)\times(M-1)$ submatrix of $B$ with determinant ±$\pm 1$. Therefore
$\operatorname{rank}(A)\;\geq\;M-1.$
Combining (i) and (ii) gives
$\operatorname{rank}(A)\;=\;M-1.$
[…]

### v3

@@ 5.2 Effect of Parameters $s$ and $t$
In GS-BCDNS, the sampling parameters $s,t\in(0,1)$ determine the working-set size $\lvert\mathcal{H}_{k}\rvert\approx(M-1)+sN$ and the number of candidate variables whose reduced costs are inspected before forming $\mathcal{I}_{k}$. This subsection investigates how $(s,t)$ affects computational cost in practice.
Test Problem 1. (Uniform–Normal, varying $s$ and $t$)
We compute the Wasserstein distance between the uniform distribution on [1,1][-1,1] and the standard normal distribution $\mathcal{N}(0,1)$ by drawing $x_{i}\sim\mathrm{Unif}[-1,1]$ and $y_{j}\sim\mathcal{N}(0,1)$ using $i.i.d.$ samples of equal size $n\in\{250,500,1000\}$.
For each pair $(s,t)$, we run GS-BCDNS until the optimality test is satisfied and record: (i) the total number of reduced-cost evaluations (denoted “reduced-cost evals”), (ii) the total number of pivots performed by the network simplex solver across all subproblems (denoted “pivot count”), and (iii) wall-clock runtime in seconds. The reduced-cost evaluations measure the dominant arithmetic workload, because each evaluation is an $O(1)$ operation in the OT network structure, and they are performed repeatedly during candidate screening and optimality checks.
Comment on Test Problem 1.
Figures 13 visualize the three metrics as heatmaps over $(s,t)$. The white regions correspond to parameter pairs that were not evaluated (in particular, those that violated $s<t$).
Across all $n$, the runtime heatmap closely matches the reduced-cost evaluation heatmap, indicating a strong positive correlation between these two quantities: configurations that trigger more reduced-cost computations consistently take longer in wall-clock time. Moreover, for all three problem sizes, both reduced-cost evaluations and runtime attain their smallest values when $s$ is chosen on the order of $2/n$ – -$3/n$ and $t$ is on the order of $20/n$ – -$30/n$, i.e.,
$s\in\Bigl[\tfrac{2}{n},\,\tfrac{3}{n}\Bigr],\qquad t\in\Bigl[\tfrac{20}{n},\,\tfrac{30}{n}\Bigr].$
suggesting that a small working-set fraction together with a moderate candidate screening fraction offers the best trade-off between (a) selecting sufficiently good entering candidates and (b) avoiding excessive reduced-cost screening overhead.
In contrast, the pivot count depends much more strongly on $s$ than on $t$. As $s$ increases, the subproblem size grows and the network simplex solver needs fewer pivots to achieve the optimality of the subproblem, leading to a clear decrease in pivot count. However, a larger $s$ also increases the per-iteration burden through a larger $\lvert\mathcal{H}_{k}\rvert$ and more bookkeeping within each subproblem; consequently, beyond a certain threshold, the reduced-cost workload increases again even though the pivot count continues to fall. This shows that the pivot count alone is not a reliable proxy for the total runtime.
In general, qualitative patterns are consistent for $n=250,500,1000$, while absolute magnitudes increase with problem size as expected.
Figure 1: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=250$.
Figure 2: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=500$.
Figure 3: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=1000$.
@@ 5.3 Comparison with the Existing Methods
In this subsection, we compare the proposed GS-BCDNS with the existing method.
First, we compare the proposed GS-BCDNS with the classical NS method and the baseline RS-BCDNS method on Test Problem 2 and 3. We also report (i) the wall-clock runtime and (ii) the total number of reduced-cost evaluations. Consistent with this, the experiments show a strong correlation between the reduced-cost evaluations and the runtime. Second, We compare GS-BCDNS with the stabilized Sinkhorn algorithm on Test Problem 4. For Sinkhorn, we test multiple regularization levels $\varepsilon\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$ under the same stabilization and stopping rule as in the previous experiments. For evaluation, we report the optimality gap $\mathrm{Gap}(t)=f(x(t))-f^{*}$ as a function of wall-clock time.
Test Problem 2 (Uniform–Normal, varying $n$).
We compute the Wasserstein distance between the uniform distribution on [1,1][-1,1] and the standard normal distribution $\mathcal{N}(0,1)$ by drawing $x_{i}\sim\mathrm{Unif}[-1,1]$ and $y_{j}\sim\mathcal{N}(0,1)$. We vary the problem size as $n\in\{50,100,150,200,250,300,350,400\}$.
Test Problem 3 (Normal–Mixture, varying $n$).
We compute the Wasserstein distance between $\mathcal{N}(0,1)$ and a two-component Gaussian mixture
$\frac{1}{2}\,\mathcal{N}(-2,1)\;+\;\frac{1}{2}\,\mathcal{N}(2,1),$
by sampling $x_{i}\sim\mathcal{N}(0,1)$ and $y_{j}\sim\frac{1}{2}\mathcal{N}(-2,1)+\frac{1}{2}\mathcal{N}(2,1)$. The same range of $n$ is tested.
Comment on Test Problem 2.
Figure 4 shows that GS-BCDNS consistently achieves the smallest runtime and the fewest reduced-cost evaluations over all tested $n$. Moreover, the advantage becomes more pronounced as $n$ increases: GS-BCDNS reduces the number of reduced-cost evaluations by roughly an order of magnitude compared with NS. In contrast, RS-BCDNS provides only modest improvements over NS. The corresponding speedup curves in Figure 5 indicate that GS-BCDNS attains large and increasing speedups with respect to NS as $n$ grows, whereas RS-BCDNS stays close to the baseline.
Comment on Test Problem 3.
The Normal–Mixture instances are empirically more challenging for random block construction. As shown in Figure 6, GS-BCDNS remains the best-performing method across all tested sizes, again substantially reducing reduced-cost evaluations and runtime. However, RS-BCDNS can become slower than NS for medium-to-large $n$. This phenomenon is consistent with the mechanism of RS-BCDNS: random augmentation tends to include many variables that are unlikely to belong to the final optimal basis, leading to repeated subproblem solves with limited progress while still incurring significant reduced-cost computations. The speedup plot in Figure 7 confirms that RS-BCDNS may fall below the NS baseline, whereas GS-BCDNS maintains a clear advantage.
Figure 4: Test Problem 2: runtime (left) and total reduced-cost evaluations (right) for NS, RS-BCDNS, and GS-BCDNS.
Figure 5: Test Problem 2: speedup relative to NS (NS time / method time).
Figure 6: Test Problem 3: runtime (left) and total reduced-cost evaluations (right) for NS, RS-BCDNS, and GS-BCDNS.
Figure 7: Test Problem 3: speedup relative to NS (NS time / method time).
Test Problem 4 (1D/2D Uniform–Normal).
Wasserstein distance estimation between (i) a one-dimensional uniform distribution on [1,1][-1,1] and a standard normal distribution, and (ii) their two-dimensional counterparts on $1,1]^{2}$ and $\mathcal{N}(0,I_{2})$.
Comment on Test Problem 4.
Figures 89 show a consistent accuracy–speed trade-off. With relatively large regularization, Sinkhorn quickly reduces the objective at the beginning, but the gap stagnates at a non-negligible level due to the inherent regularization bias. As $\varepsilon$ decreases, Sinkhorn becomes progressively closer to the unregularized OT solution, yet its convergence speed deteriorates substantially. In contrast, GS-BCDNS exhibits a slower initial decrease than low-precision Sinkhorn, but it continues to make steady progress and ultimately attains the exact optimal solution of the original OT problem.
Figure 8: Test Problem 4 (1D): optimality gap versus time for GS-BCDNS and stabilized Sinkhorn under different regularization levels $\varepsilon$.
Figure 9: Test Problem 4 (2D): optimality gap versus time for GS-BCDNS and stabilized Sinkhorn under different regularization levels $\varepsilon$.
5.4 Large-scale Experiment
This subsection evaluates the proposed method on a larger OT instance with $n=4000$, where solving the full problem directly becomes significantly more demanding. We consider a one-dimensional Wasserstein distance problem between a uniform source distribution and a heavy-tailed target distribution.
Test Problem 5 (Uniform–Beta).
The source samples $\{u_{i}\}_{i=1}^{n}$ are drawn from $\mathrm{Unif}([-1,1])$. The target samples $\{v_{j}\}_{j=1}^{n}$ are drawn from a $\mathrm{Beta}(0.5,0.5)$ distribution and then linearly mapped to [1,1][-1,1].
To interpret the evolution of the computed transport plan, we use the barycentric projection
$T_{k}(u_{i})\;:=\;\frac{1}{p_{i}}\sum_{j=1}^{n}v_{j}\,x^{k}_{ij},\qquad i=1,\dots,n,$
where $x^{k}$ denotes the transport plan produced after $k$ iterations. Intuitively, $T_{k}(u_{i})$ is the average target location to which the mass in $u_{i}$ is transported under $x^{k}$.
Comment on Test Problem 5.
Figure 10 summarizes the large-scale experiment. The top panel plots the empirical marginals: the source is approximately uniform, while the target (Beta(0.5,0.5)(0.5,0.5) mapped to [1,1][-1,1]) concentrates more mass near the two endpoints, producing a characteristic U-shaped density. This structural mismatch implies that the optimal coupling must transport a nontrivial portion of the mass from the middle region of the uniform source toward the two extremes of the target support.
The bottom panel shows the barycentric projection curves $T_{k}(x)$ at several epochs ($k=10,20,50,100$) together with the final optimal map (“Optimal”). At early epochs (e.g. $k=10$ and $k=20$), the barycentric projection exhibits substantial oscillations, reflecting that the current sparse basic solution still contains non-optimal variables and the coupling has not yet stabilized. As the iterations proceed, the projection becomes increasingly structured and approaches a smooth monotone map: by $k=50$ the curve already aligns closely with the optimal trend, and by $k=100$ it nearly overlaps the optimal map over most of the domain.
Consistent with this visual evidence, the proposed method reaches the globally optimal solution at $218$ epochs in this experiment. In general, Figure 10 demonstrates that even at $n=4000$, the iterations produced by GS-BCDNS steadily reduce the optimality gap and the associated transport map progressively converges toward the exact optimal coupling.
Figure 10: Large-scale OT experiment with $n=4000$. Top: empirical source and target marginals (Uniform[1,1][-1,1] vs. Beta(0.5,0.5)(0.5,0.5) mapped to [1,1][-1,1]). Bottom: barycentric projection $T_{k}(x)$ at selected epochs $k$, illustrating progressive convergence of the transport map toward the optimal monotone coupling.
@@ 6.  Conclusion
In this paper, we studied efficient algorithms to solve the OT problem. After reviewing representative approaches, including the NS method, BCD schemes, and the Sinkhorn algorithm, we proposed the BCDNS method. The proposed framework integrates NS pivots with block coordinate updates and introduces the basis variable succession strategy to preserve feasibility across subproblems. We established theoretical guarantees showing that the BCDNS method terminates in a finite number of iterations and converges to an exact optimal solution.
Extensive numerical experiments demonstrate that the BCDNS method substantially improves computational efficiency over the classical NS method on large-scale instances, achieving speed-ups of up to tens of times while maintaining exact optimality. Compared with the Sinkhorn algorithm, the BCDNS method exhibits a different accuracy–efficiency trade-off: although it converges more slowly than Sinkhorn with large regularization parameters, it avoids regularization bias and guarantees convergence to the OT optimum. Moreover, when high accuracy is required, the BCDNS method can be competitive with or even faster than Sinkhorn. These results indicate that the BCDNS method is particularly well suited for large-scale OT problems where exact solutions are required.
Several directions merit further investigation. First, more sophisticated block selection strategies may further accelerate convergence. Second, extending the proposed framework to more general network flow and structured linear programming problems could broaden its applicability.